OpenStudy (anonymous):

factor completely 1 - w8

OpenStudy (saifoo.khan):

1 ( 1-w8)

myininaya (myininaya):

1-(w^4)^2=(1-w^4)(1+w^4)=(1-(w^2)^2)(1+w^4)=(1-w^2)(1+w^2)(1+w^4) =(1-w)(1+w)(1+w^2)(1+w^4)

OpenStudy (anonymous):

myininaya these are the choices I have to choose from: a. (1 + w4) (1 + w2) (1 - w2) b. (1 + w4) (1 + w2) (1 + w) (1 - w) c. (1 + w4) (1 - w4) d. not factorable

OpenStudy (anonymous):

myininaya these are the choices I have to choose from: a. (1 + w4) (1 + w2) (1 - w2) b. (1 + w4) (1 + w2) (1 + w) (1 - w) c. (1 + w4) (1 - w4) d. not factorable

myininaya (myininaya):

ok whats wrong?

myininaya (myininaya):

you do know multiplication is commutative right? examples: a*b*c=c*a*b=b*a*c and so on.. 3*2=2*3 2*4*3=4*2*3 and so on...